Cathy found a coin purse which contains eight more 10-peso coin than 5-peso coin. If there are 29 coins in all, how much money is in the coin purse?
Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

5 PESO COINS=X
10 PESO COINS=(X+8)
X+(X+8)=29
X+X+8=29
2X=29-8
2X=21
X=21/2
X=10.5 5 PESO COINS.
THIS ISN'T A VALID ANSWER.
HAVE YOU MISSTATED THIS PROBLEM??????

Answer by stanbon(57424)

Cathy found a coin purse which contains eight more 10-peso coin than 5-peso coin. If there are 29 coins in all, how much money is in the coin purse?
------------
Let # of 10-peso coins be "x" ; Then # of 5-peso coins is "x-8".
EQUATION:
x + x-8 = 29
2x = 37
x = 37/2
--------------
Comment: That value is not possible for the number of 10-peso coins as the
number must be a whole number.
--------------
Check your problem statement.
=============
Cheers,
Stan H.